hello.
Please help me.
w - 5/2 - w + 2/3 = -2
I am multiplying both sides by 6. I am getting -1. However, when I plug -1 in for w, the sides are not equal.
thank you

Answer by tutorcecilia(2144)

6(w - 5/2 - w + 2/3 = -2)
6w-15-6w+4=-12
-11=-12
-11 does not equal -12, therefore, there is no solution.